On June 30, 2021, Georgia-Atlantic, Inc. leased warehouse equipment from IC Leasing Corporation. The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $530,475 over a four-year lease term (also the asset’s useful life), payable each June 30 and December 31, with the first payment at June 30, 2021. Georgia-Atlantic's incremental borrowing rate is 10%, the same rate IC used to calculate lease payment amounts. IC purchased the equipment from Builders, Inc. at a cost of $3.6 million.

Required:
1. What pretax amount related to the lease would IC report in its balance sheet at December 31, 2021?
2. What pretax amount related to the lease would IC report in its income statement for the year ended December 31, 2021?
Answer #1


Solution 1:
Semiannual lease payment = $530,475
Total semiannual payments = 4*2 = 8
Incremental borrowing rate = 10%, 5% semiannual
Present value of minimum lease payments = Semi Annual lease payments * Cumulative PV Factor of annuity due for 8 periods at 5%
= $530475 * 6.78637 = $3,600,000

Semiannual payment on 30.06.2021 = $530,475
Amount receivable on 30.06.2021 = ($3,60,000 - $530475) = $3,069,525
Interest revenue for 31.12.2021 = $3,069,525 * 5% = $153,476
Semiannual lease payment on 31.12.2021 = $530475
Pre tax amount of net receivables IC report in its balance sheet at December 31, 2021 = $3,069,525 + $153476 - $530475 = $2,692,526


Solution 2:
Pre tax amount of interest revenue IC report in its income statemetn for year ended Dec 31, 2021 = $3,069,525 * 5% = $153,476